1. A traveling wave in a string is given by where y and x are in meters and t in seconds. Find the amplitude, wavelength, the frequency, the period and the speed of the wave.
1
Expert's answer
2020-11-03T10:32:15-0500

A transverse wave is described by the following equation:


"y(x,t)=A\\text{ sin}\u2061(kx-\\omega t)."

Amplitude: "A"

Wavelength:


"\lambda=\frac{2\pi}{k}."

Frequency:


"f=\frac{\omega}{2\pi}."

Period:


"T=\frac{1}{f}=\frac{2\pi}{\omega}."

Speed:


"v=\frac{\omega}{k}."